OpenStudy (anonymous):

A 14.5 gram sample of gas is found to have a volume of 2.8 liters at 275 K and 1.10 atm. What is the molar mass of this gas? Show all of the work used to solve this problem.

OpenStudy (anonymous):

@AccessDenied

OpenStudy (anonymous):

please help I have a few like this i just need a formula

OpenStudy (accessdenied):

Have you tried Ideal Gas Law here? You can put the information about grams of gas aside and find the moles of gas, then convert using the same "mass / molar mass = moles" formula. Both of these formulas were featured in the last problem. :)

OpenStudy (anonymous):

PV=nRT??

OpenStudy (anonymous):

Pn=VRT n=VRT/P n=(14.5)(0.0821)(275)/(1.10) n=297.61

OpenStudy (anonymous):

i have this so far but I dont think it's right

OpenStudy (accessdenied):

What happened between: PV = n RT and Pn = V RT

OpenStudy (anonymous):

I really dont know I found it on yahoo and a few other places but now I'm using it it looks wrong

OpenStudy (accessdenied):

Alright. This will be an Algebra step. Note that everything here is being multiplied right now: P * V = n * R * T. To solve for n, we have to use inverse operations, so multiplying's inverse is dividing, to solve for n. And an equation is only correct if we do the same operation on both sides. Keep those two details in mind. There's an R multiplied next to n. So we can divide both sides by R. And there's also the T multiplied, so we divide that as well. The whole step is just dividing (R*T) off from both sides: \( \displaystyle \frac{P \times V}{R \times T} = \frac{n \times R \times T}{R \times T} \) Things in both numerator and denominator cancel, because anything divided by itself has to be 1. \( \displaystyle \frac{P \times V}{R \times T} = \frac{n \cancel{\times R \times T}}{\cancel{R \times T}} \) So in the end we have: \( \displaystyle \dfrac{PV}{RT} = n \) Does anything here seem unclear? The next step will be to plug in all the information!

OpenStudy (anonymous):

So (1.10)(24.5) / (0.0821)(275) = n 26.95 / 22.5775 = n 1.1937 = n

OpenStudy (accessdenied):

Looks good! Now that we have moles of gas under these conditions, and we know the mass (grams) of the gas as well, we just use the mass / molar mass = moles of gas formula. \( \dfrac{m}{M} = n \) We just have to multiply both sides by M (molar mass) to get it back on top. \( m = n \times M \) Then divide both sides by n (moles): \( \dfrac{m}{n} = M \) If everything looks good here, you have the mass (m) and moles (n) you just found, so you plug in one more time!

OpenStudy (accessdenied):

Ohh, hang on. (1.10)((((24.5))) / (0.0821)(275) = n This 24.5 looks like a typo. Volume was 2.8 liters in the problem. The other values were correct.

OpenStudy (accessdenied):

With this change, it seems you get: n = ((1.10)*(2.8))/((0.0821)*(275)) = 0.1364 mol Apologies for not catching that typo earlier!

OpenStudy (anonymous):

Sorry i left for a second lol so where do you go from there?? Or is that the answer?

OpenStudy (anonymous):

And I was looking at another problem at the same time thats probably where the 24.65 came from sorry

OpenStudy (accessdenied):

That was the corrected value of n. I didn't catch in your earlier response that you had V = 24.5, when V should be 2.8. So from here, we'd have to use the formula for mass / molar mass = moles. Oh, no problem.

OpenStudy (anonymous):

So its be molar mass = moles x mass?

OpenStudy (anonymous):

itd

OpenStudy (accessdenied):

not quite, mass divided by moles. The units make sense then: Molar mass (in grams per mole) = mass (grams) divided by moles (mol).

OpenStudy (anonymous):

how do you find the mass if the gas isn't named?

OpenStudy (anonymous):

sorry

OpenStudy (accessdenied):

That part was given: "A 14.5 gram sample of gas ..." And we have moles = 0.1364 mol, that was found a bit ago. Just to make sure it was corrected. So molar mass = mass (14.5 g) / moles (0.1364 mol).

OpenStudy (anonymous):

so molar mass =106.3049

OpenStudy (accessdenied):

Alright, that looks good now. :) (Originally I checked the end result with the moles = 1.1 to be about 12 g/mol, which was carbon... but gaseous carbon is unheard of at 275 K (room temperature...)

OpenStudy (anonymous):

I know it's ridiculous haha

OpenStudy (accessdenied):

Aye, so everything makes sense there? Anything else you might need help with? :)

OpenStudy (anonymous):

Would I use the same formulas and everything for :A sample of gas at 298 Kelvin and 0.98 atmospheres has a density of 1.45 g/L, what is the molar mass of this gas? Show all of the work used to solve this problem.

OpenStudy (accessdenied):

It will be the same formula, except some more tricky work. Density is mass per volume, and PV = n RT only has the volume. But we do know that other relationship of n (moles) = mass (grams) / molar mass (grams/mol). Let's make that substitution: \( PV = n RT \) \( n = \dfrac{m}{M} \) \( P\times V = \dfrac{m}{\color{red}{\mathbf{M}}}\times R \times T \) Because we want molar mass (highlighted red now), we need to solve for that. What should end up happening is that we'll have \(\dfrac{m}{V} \) somewhere, which is the same as density.

OpenStudy (accessdenied):

So the steps from that point would just be, multiply both sides by M (to get it on top). Then divide both sides by P*V, which will isolate M. I left out the details now so I don't have 400 lines, but I can explain it if needed. You'll get to this point: \( M = \color{green}{\dfrac{m}{V}} \times \dfrac{R\times T}{P \times V} \) And now we just replace m / V by density. Usually density is written as a greek letter, rho \(\rho\), but most people will cann it p anyways. \( M = \mathbf{\color{green}\rho} \times \dfrac{R \times T}{P \times V} \)

OpenStudy (accessdenied):

call*, not cann. lol

OpenStudy (anonymous):

Okay give me a minute to do this lol

OpenStudy (accessdenied):

Brain fart, one thing: \( M = \color{green}{\dfrac{m}{V}} \times \dfrac{R\times T}{P} \) \( M = \color{green}{\rho} \times \dfrac{R\times T}{P} \) I moved V underneath m, but then I made another one next to P. There wasn't really anything to substitute for V anyways. I'm not doing very well with typos today. o.o

OpenStudy (anonymous):

okay i was just about to ask lol dont worry about it

OpenStudy (anonymous):

I'm stuck on plugging things in

OpenStudy (anonymous):

nevermind brain fart

OpenStudy (anonymous):

afk sorry and thank you so much

OpenStudy (anonymous):

I havePV=(m/M)RT M=p(RT/P) M=1.45(0.98)(298)/(0.0821) M=1.45(292.04)/(0.0821) M=423.458/(0.0821) M=5,157.8319

OpenStudy (accessdenied):

Looks like you just flipped around R and P. R = 0.0821, and P = 0.98